### v1

# The Astrometric Resoeccentric Degeneracy: Eccentric Single Planets Mimic 2:1 Resonant Planet Pairs in Astrometry

Abstract
Detections of long-period giant exoplanets will expand dramatically with Gaia Data Release 4 (DR4), but interpreting these signals will require care. We derive the astrometric resoeccentric degeneracy: an astrometric analogue of the well-known radial velocity degeneracy in which a single eccentric planet can mimic two circular planets near a 2:1 period ratio. To first order in eccentricity, the sky-projected motion of a single eccentric orbit decomposes into a fundamental mode and first harmonic with an amplitude proportional to that eccentricity. A pair of coplanar, circular planets in a 2:1 orbital resonance produces the same harmonic structure: the outer planet sets the fundamental mode, while the inner planet supplies an apparent first harmonic. We present a mapping between the harmonic amplitudes and effective eccentricity ($e_{\mathrm{eff}}$) of a single planet that mimics a 2:1 configuration, demonstrating that $e_{\mathrm{eff}}=\,2^{1/3}(M_{p,2}/M_{p,1})$, the masses of the inner and outer planets, respectively. Using simulated Gaia data we show that (1) coplanar 2:1 systems are statistically indistinguishable from a single eccentric planet and (2) mutual inclination can break this degeneracy. This bias favors detecting mutually inclined systems, often fingerprints of a dynamically hot history – traces for processes such as planet-planet scattering or secular chaos. Determining the planetary architectures in which this degeneracy holds will be essential for measuring cool-giant occurrence rates with Gaia and for inferring their dynamical evolution histories.
software: This work made use of the following software packages: matplotlib (Hunter:2007), numpy (numpy), python (python), JAX (jax2018github), jaxoplanet (jaxoplanet), and gaiascanlaw. Software citation information aggregated using The Software Citation Station (software-citation-station-paper; software-citation-station-zenodo).
1 Introduction
Astrometric detection of exoplanets is of particular interest at the moment as we approach the era of Gaia epoch astrometry (Gaia2016). Gaia Data Release 4 (DR4) is scheduled to release the first $\sim$5.5 years of time series astrometry for up to $\sim$2.7 billion sources across the complete sky in December 2026.111According to https://www.cosmos.esa.int/web/gaia/data-release-4 accessed on November 30, 2025. Gaia’s ability to search for giant exoplanets on distant orbits via the reflex motion they induce on their host stars has long been appreciated (Bernstein1995; Casertano1996; Lattanzi2000; Casertano2008; Perryman2014; Sozzetti2014; Ranalli2018). Recently, lammers2025exoplanet re-investigated the exoplanet yield from Gaia DR4 and the subsequent mission-length $\sim$10.5 years of time series astrometry in Gaia DR5. This analysis included improvements in our understanding of distant giant-planet occurrence, the local stellar population, and Gaia astrometric precision. Via injection recovery tests, lammers2025exoplanet demonstrated that Gaia will be most sensitive to super-Jupiters ($3-13\,M_{\rm Jup}$) on $2-5\,\mathrm{AU}$ orbits around GKM-type stars ($0.4-1.3\,M_{\odot}$) within $500\,\mathrm{pc}$, predicting $\sim$1,900 planets in DR4 and $\sim$38,000 planets in DR5 with masses and orbital periods determined to better than 20%. However, they assumed an underlying population solely composed of single-planet systems, which, of course, will not be the true underlying population. An important scientific result from Gaia will be precise constraints on cool gas giant occurrence rates and so it is essential to understand degeneracies and multi-modalities that will be present in the Gaia DR4 and DR5 datasets in order to not bias these occurrence rate determinations.
Radial-velocity (RV) surveys are known to suffer from a degeneracy in which a system of two planets on nearly circular orbits in or near a 2:1 mean-motion resonance can be well fit by a single-planet model with a moderately eccentric orbit – henceforth called the resoeccentric degeneracy. This effect was first analyzed in detail by AngladaEscude2010, where it was shown that $\sim$35% of published single-planet eccentric solutions are statistically compatible with circular two-planet 2:1 configurations. In the RV case, the Keplerian signal of a single eccentric orbit can be expanded as a sum of a fundamental mode at frequency $n$ and a first harmonic at frequency $2n$ – with the harmonic amplitude $\simeq e$ times the fundamental. A two-planet circular 2:1 configuration naturally reproduces the same harmonic structure, as the outer planet mimics the fundamental frequency and the inner planet mimics the first harmonic.
Here, we derive the astrometric analogue of this effect. We show that, to first order in eccentricity, a single eccentric astrometric orbit also decomposes into a fundamental mode at $n$ and a first harmonic at $2n$ in each sky coordinate. A coplanar, circular 2:1 system produces the same harmonic content, so its astrometric signal lies in the same functional space as that of a single eccentric orbit. Mutual inclination introduces additional degrees of freedom that are not representable by a single orbit and therefore will break the degeneracy in sufficiently precise astrometric data.
The connection between this degeneracy and astrometric observations was predicted in AngladaEscude2010 – where it was noted that astrometry could provide ways to uncover eccentric imposters for systems in non-coplanar orientations where the inherent degeneracies will be broken. However, the analytics were not explicitly presented for the astrometric case; rather, AngladaEscude2010 points the reader to analyses of frequency decomposition for astrometric detection in Konacki2002.
We have organized this paper as follows. In §2 we derive the astrometric degeneracy between circular coplanar 2:1 resonant systems and single eccentric orbits via fundamental and first harmonic frequency analysis (good to first order in eccentricity). In §3 we demonstrate this degeneracy via simulated Gaia DR4 ($\sim$5.5 years) and DR5 ($\sim$10.5 years) data. In §4 we discuss the implications of this degeneracy in astrometric datasets, such as Gaia, and its impact on interpreting the demographics that arise from Gaia as well as the dynamical implications of potential detections of mutually inclined 2:1 resonant planets.
2 The Resoeccentric Degeneracy
Here we derive the astrometric resoeccentric degeneracy. In this derivation, we adopt the orbital coordinate frame as described in Feng2024 and assume purely Keplerian orbits.
The apparent astrometric displacement can be described in terms of the scaled Thiele-Innes constants $A^{\prime},B^{\prime},F^{\prime},G^{\prime}$ (Thiele1883), defined as
$A^{\prime}$ $=$ $\cos\Omega\cos\omega-\sin\Omega\sin\omega\cos i,$ (1)
$B^{\prime}$ $=$ $\sin\Omega\cos\omega+\cos\Omega\sin\omega\cos i,$ (2)
$F^{\prime}$ $=$ $-\cos\Omega\sin\omega-\sin\Omega\cos\omega\cos i,$ (3)
$G^{\prime}$ $=$ $-\sin\Omega\sin\omega+\cos\Omega\cos\omega\cos i,$ (4)
where $\Omega,\omega,i$ are the familiar orbital elements: $\Omega$ the longitude of ascending node, $\omega$ the argument of periastron of the star, and $i$ the inclination of the planet.
The stellar reflex motion projected onto the sky plane of a star induced by a single companion planet can be written, in RA $(\Delta\alpha^{\ast}=\Delta\alpha\cos\delta)$ and declination $(\Delta\delta)$ space, as
$\Delta\alpha^{\ast}(t)=\frac{Y(t)}{d}=\frac{B^{\prime}\,x(t)+G^{\prime}\,y(t)}{d},$ (5)
$\Delta\delta(t)=\frac{X(t)}{d}=\frac{A^{\prime}\,x(t)+F^{\prime}\,y(t)}{d},$ (6)
where $d$ is the distance to the star. Here, $x(t)$ and $y(t)$ are coordinates in the orbital plane, before any projection onto the sky. $X(t)$ and $Y(t)$ are the projected sky-plane coordinates of the star’s reflex motion (in the local tangent-plane frame, before scaling by distance).
2.1 Single mildly eccentric orbit: harmonic structure
In the orbital plane, the coordinates of the stellar reflex motion due to a single planet are
$x(t)$ $=$ $a_{\star}(\cos E(t)-e),$
$x(t)$ $\approx$ ${}a_{\star}\left[\cos M-\frac{3}{2}e+\frac{1}{2}e\cos 2M\right],$ (7)
$y(t)$ $=$ $a_{\star}\sqrt{1-e^{2}}\,\sin E(t),$
$y(t)$ $\approx$ ${}a_{\star}\left[\sin M+\frac{1}{2}e\sin 2M\right],$ (8)
where we have solved Kepler’s equation to first order in $e$. Here, $a_{\star}$ is the stellar reflex semi-major axis ($a_{\star}=\frac{M_{p}}{M_{p}+M_{\star}}\,a$), with $M_{p}$ and $M_{\star}$ the planet and stellar masses, respectively, and $a$ the semi-major axis of the planet with respect to the star. $E(t)$ is the eccentric anomaly, $e$ is the eccentricity, $M\equiv nt$ is the mean anomaly and $n=2\pi/P$ is the mean motion. Plugging into Equation 5 and Equation 6,
$\Delta\alpha^{\ast}(t)$ $\approx$ $\frac{a_{\star}}{d}\!\left[B^{\prime}\cos M+G^{\prime}\sin M\right]$ (9)
${}+\,\frac{1}{2}e\,\frac{a_{\star}}{d}\!\left[B^{\prime}\cos 2M+G^{\prime}\sin 2M\right]$
${}-\,\frac{3}{2}e\,\frac{a_{\star}}{d}B^{\prime},$
$\Delta\delta(t)$ $\approx$ $\frac{a_{\star}}{d}\!\left[A^{\prime}\cos M+F^{\prime}\sin M\right]$ (10)
${}+\,\frac{1}{2}e\,\frac{a_{\star}}{d}\!\left[A^{\prime}\cos 2M+F^{\prime}\sin 2M\right]$
${}-\,\frac{3}{2}e\,\frac{a_{\star}}{d}A^{\prime}.$
Each coordinate of reflex motion thus consists of:
  1. 1.
    a fundamental mode at frequency $n$,
  2. 2.
    a first harmonic at $2n$,
  3. 3.
    a constant (time-invariant) term,
with the harmonic amplitudes proportional to $e$. The constant terms represent a shift of the apparent orbit center and are degenerate with the fitted mean position; the time-dependent part of the signal is therefore fully described, to first order, by the fundamental mode and first harmonic. For our purposes, these constant terms can be safely discarded.
2.2 Two circular planets in a 2:1 coplanar configuration
Now consider two planets on circular orbits around the same star with periods $P_{1}=2P_{2}$ and stellar reflex semi-major axes $a_{\star,1}$ and $a_{\star,2}$. We assume that the orbits are coplanar and share the same $(\Omega,i,\omega)$, so that the same scaled Thiele-Innes constants $(A^{\prime},B^{\prime},F^{\prime},G^{\prime})$ apply to both planets. Assuming the outer planet shares an orbital period with the single planet analyzed in §2.1 so $P_{1}=P$, in the orbital plane the total displacement may be written as a linear combination of contributions from each planet to zeroth order in eccentricity, as we assume circular orbits, or $e_{1}=e_{2}=0$:
$x_{\rm tot}(t)$ $\approx a_{\star,1}\cos M+a_{\star,2}\cos 2M,$ (11)
$y_{\rm tot}(t)$ $\approx a_{\star,1}\sin M+a_{\star,2}\sin 2M.$ (12)
Applying the Thiele-Innes relations to the combined motion, we obtain
$\Delta\alpha^{\ast}_{\rm tot}(t)$ $\approx$ $\frac{a_{\star,1}}{d}\!\left[B^{\prime}\cos M+G^{\prime}\sin M\right]$ (13)
${}+\,\frac{a_{\star,2}}{d}\!\left[B^{\prime}\cos 2M+G^{\prime}\sin 2M\right],$
$\Delta\delta_{\rm tot}(t)$ $\approx$ $\frac{a_{\star,1}}{d}\!\left[A^{\prime}\cos M+F^{\prime}\sin M\right]$ (14)
${}+\,\frac{a_{\star,2}}{d}\!\left[A^{\prime}\cos 2M+F^{\prime}\sin 2M\right].$
Thus, just as in the single-eccentric case, each coordinate consists of:
  1. 1.
    a fundamental mode at frequency $n$,
  2. 2.
    a first harmonic at $2n$.
There is no time-invariant offset here because we have not included an explicit mean position term; in practice such a term is always fitted separately and does not affect the comparison of the time-dependent harmonics.
We note that systems in exact mean motion resonant configurations will be significantly influenced by N-body interactions, causing additional effects that may differentiate the signals from Keplerian orbits (beague2013resonance; batygin2013analytic). For systems in near 2:1 mean motion resonance – a common detected configuration in exoplanet architectures (e.g., Fabrycky2014) – this Keplerian approximation should hold (MurrayDermott1999).
2.3 Effective eccentricity mapping
We now make the equivalence between the two descriptions explicitly by comparing the amplitudes of the fundamental and first harmonic terms. This yields the relations
$a_{\star,1}$ $=$ $a_{\star}$ (15)
$a_{\star,2}$ $=$ $\frac{1}{2}e_{\rm eff}\,a_{\star},$ (16)
where $e_{\rm eff}$ is the effective eccentricity of the single-planet model. Solving for $e_{\rm eff}$,
$e_{\rm eff}=2\,\frac{a_{\star,2}}{a_{\star,1}}.$ (17)
In other words, the ratio of the astrometric semi-major axes induced by the inner and outer planets sets the effective eccentricity of the single-planet orbit that reproduces the same time-dependent harmonic content (up to first order in $e$ and ignoring the constant offset). Because the Thiele-Innes constants for the two planets are identical in the coplanar, aligned case, this mapping holds simultaneously in both coordinates.
Thus, a coplanar, phase-aligned circular 2:1 configuration produces, to first order, the same astrometric harmonic content as a single eccentric orbit at the outer period, with an effective eccentricity given by Equation 17.
2.4 From harmonic amplitudes to effective eccentricity
The astrometric resoeccentric degeneracy can be related to physical quantities of interest by converting the effective eccentricity $e_{\rm eff}$ into a relation involving the planet masses and periods. The astrometric semi-major axis induced by planet $k$ for a circular orbit is (Sozzetti2005)
$a_{\star,k}=\frac{M_{p,k}}{M_{\star}}\,\frac{a_{k}}{d},$ (18)
where $M_{p,k}$ is the planet mass, $a_{k}$ is the orbital semi-major axis, $M_{\star}$ is the stellar mass, and $d$ is the distance from the observer to the system. Using Kepler’s third law, the ratio of the two astrometric amplitudes is
$\frac{a_{\star,2}}{a_{\star,1}}=\frac{M_{p,2}}{M_{p,1}}\left(\frac{P_{2}}{P_{1}}\right)^{2/3}.$ (19)
Combining equations (17) and (19) gives
$e_{\rm eff}=2\,\frac{M_{p,2}}{M_{p,1}}\left(\frac{P_{2}}{P_{1}}\right)^{2/3}.$ (20)
Equation (20) shows that the apparent eccentricity of a single-planet astrometric fit is not set solely by the inner planet’s mass relative to the outer planet, but also by the geometric lever arm provided by the period (and therefore semi-major axis) ratio. For a strict 2:1 configuration with $P_{2}=P_{1}/2$, we have
$\frac{M_{p,2}}{M_{p,1}}=2^{-1/3}\,e_{\rm eff}\simeq 0.79\,e_{\rm eff}.$ (21)
For instance, in this exact 2:1 case, an apparent eccentricity of $e_{\rm eff}=0.30$ corresponds to $M_{p,2}/M_{p,1}\approx 0.24$, meaning that an inner planet with only one quarter of the mass of the outer planet can mimic the harmonic signature of a single planet with $e\approx 0.3$. Thus, the astrometric resoeccentric degeneracy maps cleanly into a mass-period relation, providing a physically intuitive explanation of how an apparently eccentric orbit can arise from two circular, phase-aligned planets in resonance.
2.5 Mutual inclination breaks the degeneracy
If the two planets have different orbital orientations in the sky plane, $(\Omega_{1},i_{1},\omega_{1})\neq(\Omega_{2},i_{2},\omega_{2})$, they generate independent sets of Thiele-Innes constants $(A^{\prime}_{1},B^{\prime}_{1},F^{\prime}_{1},G^{\prime}_{1})$ and $(A^{\prime}_{2},B^{\prime}_{2},F^{\prime}_{2},G^{\prime}_{2})$. The sky-plane motion then takes the form
$\Delta\alpha^{\ast}$ $=$ $\frac{B^{\prime}_{1}x_{1}(t)+G^{\prime}_{1}y_{1}(t)+B^{\prime}_{2}x_{2}(t)+G^{\prime}_{2}y_{2}(t)}{d},$ (22)
$\Delta\delta$ $=$ $\frac{A^{\prime}_{1}x_{1}(t)+F^{\prime}_{1}y_{1}(t)+A^{\prime}_{2}x_{2}(t)+F^{\prime}_{2}y_{2}(t)}{d},$ (23)
which is a sum of two ellipses with, in general, different orientations. A single Keplerian orbit possesses only one set of $(A^{\prime},B^{\prime},F^{\prime},G^{\prime})$, and therefore enforces fixed relationships among the harmonic coefficients in $\Delta\alpha^{\ast}$ and $\Delta\delta(t)$. Once the two planetary contributions have different projection geometries, those relationships no longer hold, and the motion cannot be reproduced by any single orbit.
For pairs of circular-orbiting planets in a 2:1 configuration with $e_{1}=e_{2}=0$, the sky-plane coordinates depend on the mean anomaly and $\omega$ only through the mean longitude $\lambda=M+\omega$. A shift in $\omega$ can therefore be absorbed into a compensating shift in $M$, and differences in $\omega$ (or in $M$) do not create different Thiele-Innes constants for circular orbits. Such phase shifts merely change the timing of the sinusoidal motion and do not alter its projected orientation.
Consequently, for circular-orbiting planets the degeneracy is broken only by differences in inclination or longitude of ascending node. Mutual inclination produces distinct Thiele-Innes constants and therefore destroys the resoeccentric degeneracy. In contrast, perfectly coplanar circular planets with the same $(i,\Omega)$ remain degenerate with a single eccentric orbit at first order, regardless of differences in $\omega$ or in orbital phase.
3 Simulations
We demonstrate how Gaia astrometric observations may have difficulties in differentiating between two-planet systems in a 2:1 mean-motion commensurability and a single eccentric planet. Using the jaxoplanet (jaxoplanet) framework, we create a Keplerian orbit for a pair of $e=0$ giant planets with masses $M_{p,1}=12\,M_{\mathrm{Jup}}$ and $M_{p,2}=(0.79)\,(e_{\rm eff}=0.3)\,({M_{p,1}})=2.84\,M_{\mathrm{Jup}}$ and with orbital periods $P_{1}$ and $P_{2}=P_{1}/2$. Following the same procedure, we additionally create a Keplerian orbit for a single eccentric planet of mass $M_{p,1}=12\,M_{\mathrm{Jup}}$, period $P_{1}$, and eccentricity $e_{\mathrm{eff}}=0.3$. We then duplicate this effort for a system with a mutual inclination of $45^{\circ}$ to demonstrate the resoeccentric degeneracy breaking capabilities of mutually inclined orbits.
[…]
Gaia does not measure instantaneous RA and Dec, but rather the along–scan (AL) coordinate $\Delta\eta$ at each field-of-view transit. For every simulated epoch $t_{k}$ with scan angle $\psi_{k}$ (computed from the full Gaia scanning law using gaiascanlaw), we project the sky-plane reflex motion into the AL direction via
$\Delta\eta_{k}^{\rm true}=\Delta\alpha^{\ast}_{k}\cos\psi_{k}+\Delta\delta_{k}\sin\psi_{k}.$
We then add Gaussian noise with standard deviation $\sigma_{\rm fov}$ to obtain
$\Delta\eta_{k}^{\rm obs}=\Delta\eta_{k}^{\rm true}+\mathcal{N}(0,\sigma_{\rm fov}).$
The corresponding noisy sky-plane measurements follow by projecting the AL noise realization $\epsilon_{k}=\Delta\eta_{k}^{\rm obs}-\Delta\eta_{k}^{\rm true}$ back onto RA and Dec:
$\Delta\alpha^{\ast}_{\rm obs,k}=\Delta\alpha^{\ast}_{\rm true,k}+\epsilon_{k}\cos\psi_{k},$
$\Delta\delta_{\rm obs,k}=\Delta\delta_{\rm true,k}+\epsilon_{k}\sin\psi_{k}.$
The marginalized RA and Dec uncertainties are therefore
$\sigma_{\alpha^{\ast},k}=\sigma_{\rm fov}\,|\cos\psi_{k}|,\qquad\sigma_{\delta,k}=\sigma_{\rm fov}\,|\sin\psi_{k}|.$
With this approach, we can generate synthetic sky–plane (RA, Dec) as well as along-scan ($\Delta\eta$) measurements for Gaia DR4 (5.5 yrs) and Gaia DR5 (10.5 yrs). The resulting sky-plane orbits, RA/Dec time series for these simulations, and along-scan astrometric measurements ($\Delta\eta)$ are shown in Figure 1 and Figure 2. These demonstrate that [1] it is possible for a single eccentric planet to mimic a coplanar, circular 2:1 resonant planetary system in both Gaia DR4 and DR5 and [2] in certain orbital & mass configurations, mutual inclination will break this degeneracy – allowing observers to statistically distinguish between these two model hypotheses. Future work should investigate for which orbital configurations precisely will this degeneracy breaking occur in Gaia DR4 and DR5.
Figure 1: Gaia DR5-like simulated orbits (vertical line shows Gaia DR4 baseline), coplanar system ($T_{\mathrm{span}}=5.5\,\mathrm{yr}$, $i_{\mathrm{mut}}=0^{\circ}$): 2:1 system with $P_{1}=5.2\,\mathrm{yr}$, $M_{p,1}=12\,M_{\mathrm{Jup}}$, $e_{1}=0.0$, $P_{2}=2.6\,\mathrm{yr}$, $M_{p,2}=2.84\,M_{\mathrm{Jup}}$, $e_{2}=0.0$, compared to a single-planet model with $P=5.2\,\mathrm{yr}$, $e=0.3$, $M_{p}=12\,M_{\mathrm{Jup}}$. Star assumed to be solar-mass at 50 pc. Here, $\Delta\alpha\cos\delta=\Delta\alpha^{\ast}$ is the RA and $\Delta\delta$ is the declination of the stellar reflex motion projected onto the sky plane of a star; $\Delta\eta$ is the Gaia along–scan (AL) coordinate; Res are the residuals between the simulated Gaia-like data with the eccentric model and the 2:1 model, respectively.
Figure 2: Gaia DR5-like simulated orbits (vertical line shows Gaia DR4 baseline), mutually inclined system ($T_{\mathrm{span}}=5.5\,\mathrm{yr}$, $i_{\mathrm{mut}}\approx 45^{\circ}$): 2:1 system with $P_{1}=5.2\,\mathrm{yr}$, $M_{p,1}=12\,M_{\mathrm{Jup}}$, $e_{1}=0.0$, $P_{2}=2.6\,\mathrm{yr}$, $M_{p,2}=2.84\,M_{\mathrm{Jup}}$, $e_{2}=0.0$, compared to a single-planet model with $P=5.2\,\mathrm{yr}$, $e=0.3$, $M_{p}=12\,M_{\mathrm{Jup}}$. Star assumed to be solar-mass at 50 pcpc. Here, $\Delta\alpha\cos\delta=\Delta\alpha^{\ast}$ is the RA and $\Delta\delta$ is the declination of the stellar reflex motion projected onto the sky plane of a star; $\Delta\eta$ is the Gaia along–scan (AL) coordinate; Res are the residuals between the simulated Gaia-like data with the eccentric model and the 2:1 model, respectively.
4 Implications
4.1 Exoplanet Astrometry in the Gaia Era
Gaia is projected to discover $\sim$1,900 long-period planets with the release of DR4 and $\sim$38,000 with DR5 (at <<20% precision; lammers2025exoplanet). Many more partial orbits and systems that cannot be as precisely constrained will also be observed (e.g., Wallace2025). While some giant planets on distant orbits are currently known, this dataset will represent our first population-level probe of the outer reaches of planetary systems and thus provides a unique opportunity to analyze the demographics and occurrence rates of long-period planets.
Our work highlights the meticulous care that must be taken in interpreting these demographics. We predict that if single-planet solutions are assumed, an overabundance of eccentric planets may be reported. As we shall elaborate upon in §4.3, if these eccentric planets are incorrectly taken at face value, our understanding of their dynamical histories shall be led astray. If this degeneracy is not resolved by mutual inclination as discussed in §2.5, we encourage community efforts to pursue follow-up campaigns and continue to develop rigorous sampling techniques to address it, as outlined in §4.6.
We speculate that if an unusual overabundance of planets is indeed detected at a given eccentricity, it may be possible to draw conclusions regarding the occurrence of planets of the corresponding mass ratio (see Equation 21). We defer detailed analysis of this to future work.
4.2 Long-Period Resonant Planets May Be Common
In multi-planet systems, interactions with the protoplanetary disk generally induce convergent inward migration (lin1986proto), making compact resonant chains the expected outcome of planetary formation (e.g. snellgrove2001disc; lee2002dynamics). Indeed, the outer Solar System is believed to have originated in such a resonant configuration (tsiganis2005; gomes2005; morbi2005). While dai2024resonance showed resonance is indeed common in young systems, this trend is not observed in more mature systems. The generally accepted picture is that resonant chains destabilized over the subsequent evolution of the system, a paradigm coined “breaking the chains” (izidoro2017chains).
[…]
4.3 Mutually Inclined Systems
The fact that mutual inclination eliminates this degeneracy (see §2.5) is fortuitous, given that the mutual inclination is also of paramount dynamical interest. Planets are expected to form in circular well-aligned orbits, as planet-disk interactions will damp out any eccentricity or inclination excursions (cresswell2007evolution). Hence, both eccentric planets and mutually inclined planets (e.g. xuan2020mutual; an2025significant; bgagliuffi2025jwst) may be signals of dynamically hot histories. However, the inferred evolutionary pathways differ for an isolated eccentric planet versus a system containing two misaligned planets. Below, we summarize the dynamical histories predicted by several mechanisms considered especially influential in sculpting exoplanetary architectures.
  • Planet-planet scattering (chatterjee_2008; juric2008dynamical; nagasawa2008scattering; ford2008; petrovich2014scattering; anderson2020insitu; marzari2025planet; lu2025hatp11) involves violent gravitational interactions induced by close encounters between pairs of planets, and can generate both significant eccentricity and inclination. raymond2008mmrs showed that while mean motion resonances are not the typical outcome of planet-planet scattering, this outcome does occur with appreciable frequency. They opine
    “Thus, scattering is likely to be responsible for systems past 1 AU with 2:1 or 3:2 resonant planets and a more massive outer planet.”
    Planet-planet scattering events involving only two planets tend to eject one of them, and rarely generate extremely high mutual inclinations in the cases where both planets survive (ford2001dynamical; nagasawa2011orbital). Hence, a mutually inclined system likely originated with three or more planets while a single eccentric planet could be generated with only two planets, if scattering is responsible.
  • Secular interactions between planets can result in both excited eccentricities and inclinations (MurrayDermott1999), and these effects are greatly enhanced near mean motion resonances (e.g. malhotra1998resonances). If instead a single eccentric planet must be explained via secular interactions, the dynamical interpretation changes – the perturbing planet must be distant enough as to be undetectable. For instance, the system may be hierarchical and an unseen distant body may induce ZLK oscillations (vonzeipel1910zlk; lidov1962evolution; kozai1962secular; naoz2016eccentric).
  • Primordial misalignment. While eccentricity is very difficult to explain without invoking some sort of dynamical mechanism (e.g. duffell2015eccentric), the same cannot be said for misalignments. Primordial misalignments can arise naturally through processes that tilt the protoplanetary disk itself (lai2008wave; kuffmeier2021misaligned), a scenario for which there is growing observational support (BiddleBowler2025). Intriguingly, however, we note that such observational efforts are only capable of probing the outer reaches of protoplanetary disks – there is complementary evidence suggesting that inner disks are preferentially aligned (radzom2024misalignment; radzom2025misalignment). This supports the hypothesis that disks may often warp into two mutually misaligned sections (e.g. nixon2013tearing; nealon2018warp), resulting in mutually inclined planets.
We note that dynamical mechanisms generally do not yield two perfectly circular orbits in a mean-motion resonance; some eccentricity is typically excited in the process. We emphasize that these are all first order approximations in eccentricity and thus will be increasingly less reliable as eccentricity grows.
4.4 Peas-in-a-Pod-like Architecture
Our constraints on the mass ratios of distant giant planets are poor, both observationally and theoretically. A tempting assumption to make would be that of relatively equal-mass planets, a trend seen in the demographics of closer-in planets and coined “peas-in-a-pod” (millholland2017peas; weiss2018peas).
If indeed the outer reaches of planetary systems largely exhibit the same intra-system mass uniformity as their close-in counterparts, the degeneracy we have pointed out will not be a concern. Inspection of Equation 21 reveals that should the two planets in question be of equal mass, a single planet masquerading as the pair would require an unphysical eccentricity of $e_{\mathrm{eff}}$ = 1.26 to mimic the astrometric signal. More generally, in systems where $M_{p,2}\gtrsim 0.79M_{p,1}$ this degeneracy will not be a concern.
4.5 Hill-stability mass limits
The preferred parameter space of Gaia-detectable planets will be super-Jupiters on 25AU2-5\,AU orbits (lammers2025exoplanet). At first glance, this narrow semimajor-axis range may seem too restrictive to host multiple massive planets on long-term stable orbits. In this section, we demonstrate that such configurations are in fact readily achievable.
The Hill stability boundary of two planets on initially coplanar circular orbits is well understood. Gladman1993 derived an analytic criterion for the orbital spacing $\Delta\equiv a_{2}-a_{1}$ demanded for Hill stability for two planets of masses $M_{p,1}$ and $M_{p,2}$ orbiting a star of mass $M_{\star}$:
$\frac{\Delta}{a_{1}}>2.4\,\left(\frac{M_{p,1}+M_{p,2}}{M_{\star}}\right)^{1/3}.$ (24)
We derive the maximum total planetary mass compatible with Hill stability for a pair of planets locked in a 2:1 period commensurability. From Kepler’s third law (kepler1609third), $\Delta\approx 0.587a_{1}$ for the 2:1 resonance. We may substitute this into Equation 24 and solve for the total planetary mass $M_{p,1}+M_{p,2}$:
$M_{p,1}+M_{p,2}\lesssim 0.0149M_{\star}.$ (25)
For a solar-mass host star, two planets with a total mass of up to $\sim 15.6\,M_{\rm Jup}$, easily detectable by Gaia, can be expected to be Hill stable around a solar mass star in a 2:1 period configuration – thus further emphasizing the importance for observers to prepare for the possibility of this degeneracy.
We note that Gladman1993 did not consider the stabilizing influence of mean motion resonances in their study. Numerous authors (e.g. wisdom1980resonance; marti2013gj876; obertas_2017; tamayo2017convergent; Tamayo20; lammers2024instability; lammers_winn_2024; lu_hip; tamayo2025unified) have demonstrated that exact mean motion resonances are associated with strong islands of stability. We present evidence in §4.2 that capture into these stable islands via convergent migration is an expected outcome of planet formation. This implies that larger planets may also be stable on long timescales in 2:1 configurations.
4.6 Other Ways to Break this Degeneracy
[…]
Combining radial velocity and astrometric observations is a powerful way to determine true planetary masses and orbital inclinations. Radial velocities measure only the line-of-sight motion and therefore yield the minimum mass, $M_{p}\sin i$, while astrometry measures the sky-plane component of the stellar reflex motion. When used together, these complementary data sets break the longstanding $\sin i$ degeneracy of radial velocity analyses (e.g., WrightHoward2009; Brandt2019; Winn2022; Yahalomi2023; Feng2024).
This RV-astrometry synergy has recently been demonstrated in practice via follow-up radial velocity monitoring to directly confirm and refine astrometric orbit solutions from Gaia DR3. In particular, RV follow-up of the Gaia-4b and Gaia-5b candidates revealed a massive planet and a brown dwarf orbiting low-mass stars, validating the astrometric orbits and providing precise dynamical masses (stefansson2025gaia). We note that even the Gaia-4b and Gaia-5b signals – with effective eccentricities $e_{\rm eff}\simeq 0.34$ and $e_{\rm eff}\simeq 0.64$, respectively, from the single-planet fits in stefansson2025gaia – could in principle be mimicked by circular 2:1 planet pairs: for Gaia-4, a pair with periods $\{571,286\}\,\mathrm{days}$ and masses $\{11.8,3.2\}\,M_{\rm Jup}$; and for Gaia-5, a pair with periods $\{359,179\}\,\mathrm{days}$ and masses $\{20.9,10.6\}\,M_{\rm Jup}$. While a detailed investigation of the resoeccentric degeneracy for these two systems is beyond the scope of this work, such an analysis would help clarify whether, and under what conditions, RV follow-up of astrometrically identified systems can break this degeneracy. Future work should additionally investigate whether RV breaking of the resoeccentric degeneracy is feasible for the types of planets we expect to find in Gaia DR4 and DR5 astrometry based on current RV capabilities and study the optimal orbital strategies for such observations.
This work predicts an additional degeneracy that may appear in the Gaia epoch astrometric data. It is already well-documented that astrometric solutions become highly degenerate when the orbital period extends beyond the observational baseline (e.g., Casertano2008). This highlights the need for the development of proper astrometric inference techniques that are designed to explore highly degenerate and multi-modal parameter spaces. Modern orbit-fitting frameworks such as The Joker (thejoker), orbitize! (orbitize1; orbitize2), OFTI (ofti), Octofitter (octofitter), and orvara (orvara) have demonstrated the importance of using samplers that remain robust in the presence of strong covariances, long-lived degeneracies, and disconnected orbital solutions. Techniques such as rejection sampling, parallel tempering, and nested sampling have been developed precisely for this purpose, providing reliable posterior exploration even when the likelihood surface contains multiple isolated modes or narrow curved manifolds. These approaches are essential for accurately capturing the full family of orbital solutions implied by astrometric and radial-velocity data, and we strongly encourage their further development and study in preparation for Gaia DR4 and DR5. It will be important to understand in which regions of parameter space these different sampling methods optimally sample the complete solution space.
We also encourage future work to investigate the signal-to-noise ratio (SNR) required to differentiate between the 2:1 and eccentric configurations. This will likely be a function of the masses, periods, and other orbital parameters (such as eccentricity and mutual inclination of the 2:1 planets), but will be a powerful result in ensuring over what regions of parameter space are statistically differentiated solutions possible. See kippingRVsnr2024 for the analogous derivation of this for RV data, which was built on the inclusion of higher order eccentricity terms as derived in Lucy2005.
Finally, it may be possible to detect this resoeccentric degeneracy on a population level by searching for non-isotropic solutions of $\omega$ (the argument of the periastron) of single eccentric planets fit to the Gaia data. As the $\omega$ term would need to be “fine-tuned” in order to mimic 2:1 astrometric signals, anisotropies in the resulting $\omega$ distribution are likely – one could potentially flag that this effect is present at the population level. Injection-recovery of 2:1 systems fit with only single eccentric solutions could potentially confirm this degeneracy-flagging potential before Gaia DR4.
5 Conclusion
In this work, we derived the astrometric resoeccentric degeneracy: an extension of the 2:1-eccentric degeneracy familiar from RV studies. To first order in eccentricity, the sky-plane motion of a star induced by a single eccentric planet decomposes into a fundamental frequency and a first harmonic with an amplitude that scales with its eccentricity; a coplanar pair of circular planets in a 2:1 period ratio induces the same harmonic structure. We provide a simple mass-eccentricity mapping that links the two descriptions, defining exactly when a two-planet system can masquerade as a single eccentric one and vice-versa.
Using Gaia-like DR4 and DR5 simulations, we showed that the reflex motion induced by a mildly eccentric single planet could mimic the reflex motion of a circular 2:1 planetary pair. We also demonstrate that mutual inclination can break this degeneracy by introducing distinct projection geometries for the two planets.
If single-planet fits are assumed by default, circular resonant pairs may be mistaken for single eccentric planets – biasing inferred eccentricity distributions, dynamical interpretations, and occurrence rates. This underscores the need for orbit-fitting approaches that can handle multi-modal astrometric likelihoods, including this resoeccentric degeneracy. Samplers capable of robustly exploring degenerate orbital solutions are essential to avoid collapsing 2:1 configurations into single eccentric solutions. A combined astrometry and RV approach appears particularly promising. Systems like Gaia-4b and Gaia-5b – where RV follow-up already exists – offer ideal real-world test beds for assessing when and how RV observations can break the 2:1-eccentric astrometric degeneracy.
Several natural extensions follow from this work. It will be important to map out the regions of parameter space in which mutual inclination or follow-up measurements reliably distinguish the two orbital architectures, and to quantify the signal-to-noise thresholds neede for Gaia to do the same. Population-level indicators, such as anisotropies in the fitted argument of periastron distribution, may provide additional diagnostic power. Injection-recovery experiments that include resonant multi-planet architectures – in addition to single-planet systems – will be key for unbiased occurrence-rate measurements.
Taken together, these efforts will help ensure that Gaia’s epoch astrometry yields not just the detection of many long-period giant planets but an accurate understanding of their occurrence rates, architectures, and the dynamical processes that shaped them.
We are very grateful to David Kipping for comments and feedback on an earlier version of this manuscript. We thank the CCA Astronomical Data Group and the Rice Research Group for helpful and illuminating conversations. D.A.Y. and T.L. are supported by Flatiron Research Fellowships at the Flatiron Institute, a division of the Simons Foundation. M.R. acknowledges support from Heising-Simons grants #2021-2802 and #2023-4478, as well as National Geographic grant EC-115062R-24 and NASA Exoplanets Research Program NNH23ZDA001N-XRP (grant No. 80NSSC24K0153). ChatGPT was utilized to improve wording at the sentence level and assist with coding inquires – last accessed in 2025 November